OpenStudy (anonymous):

give all the solutions of the nonlinear system of euqations, including those with nonreal complex components:

OpenStudy (anonymous):

2x+|y|=4 x^2 + y^2 = 5

OpenStudy (anonymous):

Can I use substitution??

OpenStudy (anonymous):

@jim_thompson

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

@Jhannybean

OpenStudy (zzr0ck3r):

I don't see why not.

OpenStudy (anonymous):

The absolute is confusing me

OpenStudy (zzr0ck3r):

2x+|y|=4 implies |y|=4-2x so y=4-2x and y=2x-4

OpenStudy (anonymous):

So now I sub which one? Both??

OpenStudy (zzr0ck3r):

x^2 + (2x-4)^2 = 5 and x^2 + (4-2x)^2 = 5

OpenStudy (zzr0ck3r):

both

OpenStudy (zzr0ck3r):

got it?

OpenStudy (anonymous):

Yes! So I get two solutions for x?

OpenStudy (zzr0ck3r):

prob more, two in each equation

OpenStudy (anonymous):

I got 5x^2-16x+11 = 0 for both

OpenStudy (anonymous):

Correct??

OpenStudy (zzr0ck3r):

seems so to me

OpenStudy (zzr0ck3r):

hmm sort of strange because this gives no complex solutions

OpenStudy (shubhamsrg):

you may also try to represent both graphically

OpenStudy (zzr0ck3r):

hmm just looking at this I see that 2, -2, and 1 are solutions. I am confused...

OpenStudy (anonymous):

What??

OpenStudy (zzr0ck3r):

lets try x^2 + y^2 = 5 so y=+-sqrt(5-x^2) so 2x+sqrt(5-x^2)=4 sqrt(5-x^2)=4-2x 5-x^2=(4-2x)^2 = 16-16x+4x^2 5x^2-16x-11=0, again the same solutions but I know 2 is a solution....

OpenStudy (anonymous):

Ok wait so...What? Lol

OpenStudy (anonymous):

I think you lost me, so is solution x= 2 , -2, 1??

OpenStudy (zzr0ck3r):

what are you doing in class? what class is this for?

OpenStudy (zzr0ck3r):

well I know that those are 3 solutions, and I don't know how to algebraically get 2 or -2

OpenStudy (anonymous):

Isn't it 5x2-16x+11??

OpenStudy (zzr0ck3r):

that equation has two solutions, namely 1 and 11/5 and I know 2 is a solution from looking at it. So we are missing something. I hope @shubhamsrg has an idea.

OpenStudy (shubhamsrg):

x=2 is not a solution

OpenStudy (zzr0ck3r):

that equation(5x2-16x+11) has two solutions, namely 1 and 11/5 and I know 2 is a solution to the original system from looking at it. So we are missing something. I hope @shubhamsrg has an idea.

OpenStudy (shubhamsrg):

x=1 is the only solution. which will give 2 values of y

OpenStudy (anonymous):

How @husbandry ??

OpenStudy (zzr0ck3r):

sorry its late....

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (zzr0ck3r):

ok so why is 11/5 not a solution.?

OpenStudy (shubhamsrg):

graphically it'll show you it should have only 2 real solution (1,2) and (1,-2) are the ones also, you may check 11/5 does not satisfy both

OpenStudy (shubhamsrg):

and late ?

OpenStudy (zzr0ck3r):

I know, but using the substitution method, we get x=11/5 when we solve the quadratic. So im just curious.

OpenStudy (anonymous):

can you graph what you are speaking of?

OpenStudy (shubhamsrg):

x=11/5 does not even fall on 2x + |y| = 4

OpenStudy (shubhamsrg):

I'll wolfram this.. and give you its graph..

OpenStudy (shubhamsrg):

http://prntscr.com/1d3j3l

OpenStudy (zzr0ck3r):

2x+|y|=4 implies |y|=4-2x so y=4-2x and y=2x-4 plugging this in we get 5x^2-16x+11 = 0 this has two solutions 1 and 11/5

OpenStudy (zzr0ck3r):

I understand It does not satisfy the first, but im confused then why are we getting it as a solution to the second equation depending on the first equation.

OpenStudy (zzr0ck3r):

I have not done much work with systems of non linear equations.

OpenStudy (shubhamsrg):

y^2 = |y|^2 and |y| = 4-2x since (4-2x)^2 = (2x-4)^2 , we get the result we are right now getting, i.e 11/5 as solution had the linear eqn been simply 2x+ y = 4 or 2x-y = 4, 11/5 would have been another correct answer .

OpenStudy (anonymous):

Me too, zz I am unfamiliar

OpenStudy (zzr0ck3r):

im just trying to figure out a way to see that without plugging in the solution to see if it works, or graphing it.

OpenStudy (shubhamsrg):

was I helpful in clearing the doubt ?

OpenStudy (zzr0ck3r):

there is no doupt, I still would like an analytic way to know that a solution is a solution without having to plug it in.

OpenStudy (anonymous):

Yes! I just can't see how to get the solutions, did you use sub or graph???

OpenStudy (zzr0ck3r):

well we get 2 solutions from 5x^2-16x+11 = 0 1,11/5 plug in 11/5 for x 2*11/5+|y|=4 so |y| = 4-2*11/5 <0 has no solutions but let x=1 and we see y=+-2

OpenStudy (zzr0ck3r):

because 2*1+|y|=4 implies |y| = 2 implies y=+-2

OpenStudy (zzr0ck3r):

do you understand how we can get the solutions?

OpenStudy (zzr0ck3r):

I would just like to know a way that skips the plug and check at the end.

OpenStudy (zzr0ck3r):

or logic....

OpenStudy (anonymous):

I got it. Can you help with just ONE more. I want to finish hw before class

OpenStudy (zzr0ck3r):

I can try:)

OpenStudy (anonymous):

Ok here or new box??

OpenStudy (zzr0ck3r):

maybe new one just in case im confused again:)